# Enhancing (quasi-)long-range order in a two-dimensional driven crystal

### v1

@@ III.2 Hyperuniformity and long-range order
To investigate the crystalline order in this system, we simulate crystalline configurations at $\phi=0.75$ which sets the lattice spacing $a$. This packing fraction corresponds to systems with long-range bond orientational order at equilibrium [109] but can give rise to a hexatic or a liquid phase in highly dissipative granular systems [110]. To avoid these scenarios, we choose a relatively high $\alpha\geq 0.95$ and always check that the crystalline phase is stable. For simulations at $T_{com}=0$, we stay away from the absorbing region by choosing $\gamma_{com}$ and $\Delta$ such that $\phi\gg\phi_{c}$, ensuring a sufficiently high energy injection at collision.
@@
We first perform a similar analysis as the one done in Ref. 32, fixing the COM by setting $T_{com}=0$. The results for three different systems are depicted in Fig. 2. Panels a), b) and c) show respectively the structure factor (and the static velocity factor in inset) as a function of $k$, the long-time behavior of the mean square displacement (MSD) for different system sizes and finally the translational correlation function as a function of the distance. The three quantities contain the same information expressed differently. It is illuminating to analyze all three together to better understand the theory given in Sec. II and verify its consistency. We note that the MSD of a particle in a finite crystal reaches a plateau at large times. The value of this plateau is of course dependent on the size of the cage made by neighboring particles but also on the collective excitations that induce vibrations on the crystal. From the Fourier transform of the displacement $\tilde{\bm{u}}_{k}(w)$, it can be shown that the correlation $\langle(\bm{u}_{n}-\bm{u}_{m})^{2}\rangle$ for an infinite system corresponds to MSD$(t\rightarrow\infty)$ for a system of size $(a|\bm{m}-\bm{n}|)^{2}$ in the large system size limit. We chose to use the MSD instead of the displacement correlation to check relations derived in the theory part involving the displacement correlation, the former being simpler to measure.
@@
We first focus on the curves ”Equilib” which correspond to an equilibrium system of pure hard disks ($\alpha=1$, $\Delta=0$ and $\gamma_{com}=0$). As expected, its structure factor is approximately flat (it has in reality an Ornstein-Zernike shape due to the non linear interactions [70]) and reaches a well defined value at small $q$. Its large time MSD grows as a unique logarithm of the particle number as predicted by Eq. (15) and thus its correlation function decreases as a power law with an exponent linked to the logarithmic decrease of the MSD as given by (16) and illustrated with the dashed lines. The system displays quasi-long-range order.
@@
We now turn to the $\Delta$ model, a system with only active collision and no global damping ($\alpha<1$, $\Delta\neq 0$ but $\gamma_{com}=0$). Its structure factor has a transient hyperuniform scaling before reaching a plateau at small $q$. The same behavior is found for the static velocity factor. This is incompatible with the modelization of the active collision as an FDT respecting bath at temperature $T_{loc}$ and damping $\gamma_{loc}\omega_{k}^{2}$. Indeed, if the active collisions were acting on a coarse grained level like such an equilibrium bath, the structure factor should have resembled the Equilib one. It can be shown that the fast temperature field is highly out-of-equilibrium and acts as a genuinely non-equilibrium bath for the slow velocity field. This effect induces the observed decay. This is however not the main concern of this paper and does not change the phenomenology observed when an additional global damping is added. Hence, the study of this peculiar bath is left for future investigations. The large time MSD of the $\Delta$ model displays two distinct logarithmic increases corresponding to the two plateaus reached at large and low $q$ by the structure factor. While the mechanism behind this double logarithmic scaling differs from the one arising due to the difference of temperature between a global and a local COM conserving bath in Eq. (15) (here we recall that there is no global bath since $\gamma_{com}=0$), the mathematics remain the same, as the low $k$ structure factor is related to the SDF and consequently the MSD, through Eq. (17). At small wavectors, we conclude that any 2D isotropic harmonic crystal displaying a hyperuniform scaling in $k\in[k_{1},k_{2}]$ and otherwise a constant structure factor, will exhibit this double logarithmic scaling on the displacement correlation; one in the region $r\gg 2\pi/k_{1}$ and one in $r\ll 2\pi/k_{2}$ with a transition region in between. $S(k\rightarrow 0)$ and $S(k\rightarrow\infty)$ will respectively be proportional to the prefactor of the long and short range logarithmic increase of the MSD. If the hyperuniformity extends to $k_{1}=0$ the plateau of the MSD eventually reaches a constant. The translational correlation function of the $\Delta$ model exhibits as well a double power law decrease, in line with the double logarithmic increase of the plateau of the MSD. The system displays quasi-long-range order as well.
@@ III.3 Arbitrary enhancement of quasi-long-range order via non-equilibrium effects
We now turn to the analysis of our full system with $T_{com}\neq 0$ and will show that the relevant temperature is, as expected from the theory, the temperature of the COM and not the temperature of the system $T$ or the temperature of the active collisions $T_{loc}$. Results for a range of temperatures $T_{com}$ of the global bath are presented in Fig. 3a). In the main figure, we show various curves of the plateau of the MSD versus the number of particles in the system $N$ for different $T_{com}$ in a semi-log scale. We ensure $N$ is sufficiently large to only observe the MSD scaling associated with the COM temperature. As expected from Eq. (15), the increase of the MSD is logarithmic and is greater the larger is $T_{com}$. Of course, we recover a true long-range order in the limit $T_{com}\rightarrow 0$ with a flat plateau of the MSD. In the inset, we extract the coefficient $\kappa$ in front of the logarithmic increase of the MSD as a function of $T_{com}$. That is:
@@ IV Conclusions and perspectives towards realistic systems
Quasi-2D vibrated granular fluids [114, 115, 116, 117, 118, 119, 38] represent realization of the coarse grained model discussed in section III . In these systems, macroscopic grains are confined in a box of size $L\times L\times h$ with $h\ll L$ the height of the box. The box is sinusoidally vibrated which imparts vertical energy to the grains. Collisions between particles transfer this vertical energy to the $xy$ components of the velocity. Moreover, tangential friction resulting from collisions with the top or bottom plates slows down the beads in the $xy$ plane. When viewed purely in the $xy$ plane, this system behaves similarly to the coarse-grained model discussed in section III. Indeed, collisions between particles inject energy into the $xy$ plane and play the same role as the $\Delta$ in Eq. (23) while the damping $\gamma_{com}$ during the free flight in the coarse-grained model is equivalent to the energy loss due to collisions with the roof and the bottom plate in the quasi-2D system. Moreover, the roughness of the two horizontal confining plates implies that each grain-plate collision introduces some randomness to the velocities (without even considering the angular momentum of the grains). Remarkably, this effect has been found to act as a small homogeneous thermal noise [42] which can be mapped into an effective global temperature $T_{com}$. The fact that this realistic setup is correctly described by the coarse grain model suggests the possibility of enhancing the quasi-long-range order by tuning the experimental parameters (driving amplitude, frequency, material properties, …) to reduce as much as possible the ratio between the effective $T_{com}$ caused by the plate roughness and the horizontal kinetic energy of the system $T$. Of course, the specific way in which the experimental parameters influence $T_{com}$ and $T$ can be very complex, but after some preliminary analysis, one might in principle be able to find the right combination of them to vary in order to arbitrarily reduce $T_{com}/T$. We look forward to applying these ideas to realistic granular quasi-2D systems in future studies.
@@ Appendix A Discrete real space conserving noises for vectorial quantities
The first term is a discrete Laplacian and the second is a discrete divergence of a rank 2 random tensor. Once the discrete differential operators are well defined, this equation can be understood as a discrete analogue of the model B field theory. Note however that the global conservation of momentum depends on boundary conditions.
On a general basis, any lattice can be understood as a graph $\mathcal{G}$ with a set of vertices $\mathcal{V}$ and edges $\mathcal{E}$ connecting them. The usual notion of continuous field living on $\mathbb{R}^{d}$ (or more generally on a manifold) is replaced by the notion of field living on vertices and mapping vertex to vectors. For example in our case $\bm{u}:\mathcal{V}\to\mathbb{R}^{2}$. The complexity of dealing with discrete calculus on lattices lies in the fact that the discrete gradient maps fields living on vertices to fields living on the edges while the discrete divergence maps fields living on the edges to fields living on the vertices [120, 121]. Hence, the Laplacian being the composition of the divergence with the gradient maps fields leaving on the vertices to fields living on the vertices. This immediately implies that the divergence in Eq. (29) is ill defined since we are applying it on a (random) field living on the vertices and not on the edges. Nonetheless, following Ref. 66, we will see that this term can be represented through a finite difference of random vertices field and be made well defined. Moreover, in order for the dynamics to be equilibrium-like, the discretization scheme of the differential operators as well as the variance of the noise must be well chosen in order to respect the FDT.
@@ Appendix B Fourier transform convention
Appendix C Computation of the displacement correlation function
@@
Appendix D Entropons in a two temperature delta correlated bath
@@
Appendix E Hybrid event-driven/time-stepped molecular dynamics
@@
Appendix F Temperature of the active collision bath

### v2

@@ III.2 Hyperuniformity and long-range order
To investigate the crystalline order in this system, we simulate crystalline configurations at $\phi=0.75$ which sets the lattice spacing $a$. This packing fraction corresponds to systems with long-range bond orientational order at equilibrium [111] but can give rise to a hexatic or a liquid phase in highly dissipative granular systems [112]. To avoid these scenarios, we choose a relatively high $\alpha\geq 0.95$ and always check that the crystalline phase is stable. For simulations at $T_{com}=0$, we stay away from the absorbing region by choosing $\gamma_{com}$ and $\Delta$ such that $\phi$ is above the critical packing fraction of the absorbing to active phase transition and hence, the system has a steady state with finite kinetic energy. Choosing a packing fraction such that $l(\phi)<{\Delta}/{\gamma}$ with $l$ the mean free path, ensures a sufficiently high energy injection at collision which prevents the system to reach the absorbing state [38].
@@
We first perform a similar analysis as the one done in Ref. 33, fixing the COM by setting $T_{com}=0$. The results for three different systems are depicted in Fig. 2. Panels a), b) and c) show respectively the structure factor (and the static velocity factor in inset) as a function of $k$, the long-time behavior of the mean square displacement (MSD) for different system sizes and finally the translational correlation function as a function of the distance. The three quantities contain the same information expressed differently. It is illuminating to analyze all three together to better understand the theory given in Sec. II and verify its consistency. We note that the MSD of a particle in a finite crystal reaches a plateau at large times. The value of this plateau is of course dependent on the size of the cage made by neighboring particles but also on the collective excitations that induce vibrations on the crystal. From the Fourier transform of the displacement $\tilde{\bm{u}}_{k}(w)$, it can be shown that the correlation $\langle(\bm{u}_{n}-\bm{u}_{m})^{2}\rangle$ for an infinite system corresponds to MSD$(t\rightarrow\infty)$ for a system of size $(a|\bm{m}-\bm{n}|)^{2}$ in the large system size limit. We chose to use the MSD instead of the displacement correlation to check relations derived in the theory part involving the displacement correlation, the former being simpler to measure.
@@
We first focus on the curves ”Equilib” which correspond to an equilibrium system of pure hard disks ($\alpha=1$, $\Delta=0$ and $\gamma_{com}=0$). As expected, its structure factor is approximately flat (it has in reality an Ornstein-Zernike shape due to the non linear interactions [71]) and reaches a well defined value at small $k$. Its large time MSD grows as a unique logarithm of the particle number as predicted by Eq. (15) and thus its correlation function decreases as a power law with an exponent linked to the logarithmic decrease of the MSD as given by (16) and illustrated with the dashed lines. The system displays quasi-long-range order.
@@
We now turn to the $\Delta$ model, a system with only active collision and no global damping ($\alpha<1$, $\Delta\neq 0$ but $\gamma_{com}=0$). Hence, the dynamics of the system is ballistic between collisions and the non-equilibrium effects arise purely from the non-equilibrium collision rule Eq. (23). The steady state of this system is the result of an interplay between energy injected at collision by $\Delta$ and dissipated by $\alpha$ as given by Eq. (24). In these conditions, the temperature of the system $T$ is a function of $\alpha$ and $\Delta^{2}$ (see Appendix G, Eq. (51)) and by definition it equilibrates to the local bath temperature $T_{loc}$. Its structure factor has a transient hyperuniform scaling before reaching a plateau at small $k$. The same behavior is found for the static velocity factor. This is incompatible with the modelization of the active collision as an FDT respecting bath at temperature $T_{loc}$ and damping $\gamma_{loc}\omega_{k}^{2}$. Indeed, if the active collisions were acting on a coarse grained level like such an equilibrium bath, the structure factor should have resembled the Equilib one. It can be shown that, in a hydrodynamic description [106], the fast temperature field is highly out-of-equilibrium and acts as a genuinely non-equilibrium bath for the slow velocity field. This effect induces the observed decay. This is however not the main concern of this paper and does not change the phenomenology observed when an additional global damping is added. Hence, the study of this peculiar bath is left for future investigations. The large time MSD of the $\Delta$ model displays two distinct logarithmic increases corresponding to the two plateaus reached at large and low-$k$ by the structure factor. While the mechanism behind this double logarithmic scaling differs from the one arising due to the difference of temperature between a global and a local COM conserving bath in Eq. (15) (here we recall that there is no global bath since $\gamma_{com}=0$), the mathematics remain the same, as the low-$k$ structure factor is related to the SDF and consequently the MSD, through Eq. (17). At small wavectors, we conclude that any 2D isotropic harmonic crystal displaying a hyperuniform scaling in $k\in[k_{1},k_{2}]$ and otherwise a constant structure factor, will exhibit this double logarithmic scaling on the displacement correlation; one in the region $r\gg 2\pi/k_{1}$ and one in $r\ll 2\pi/k_{2}$ with a transition region in between. $S(k\rightarrow 0)$ and $S(k\rightarrow\infty)$ will respectively be proportional to the prefactor of the long and short range logarithmic increase of the MSD. If the hyperuniformity extends to $k_{1}=0$ the plateau of the MSD eventually reaches a constant. The translational correlation function of the $\Delta$ model exhibits as well a double power law decrease, in line with the double logarithmic increase of the plateau of the MSD. The system displays quasi-long-range order as well.
@@ III.3 Arbitrary enhancement of quasi-long-range order via non-equilibrium effects
We now turn to the analysis of our full system with $T_{com}\neq 0$ and will show that the relevant temperature is, as expected from the theory, the temperature of the COM and not the temperature of the system $T$ or the temperature of the active collisions $T_{loc}$. Results for a range of temperatures $T_{com}$ of the global bath are presented in Fig. 3a). In the main figure, we show various curves of the plateau of the MSD versus the number of particles in the system $N$ for different $T_{com}$ in a semi-log scale. We also provide results for the structure factor and translational pair correlation function in the presence of global bath in Appendix H. We ensure $N$ is sufficiently large to only observe the MSD scaling associated with the COM temperature. As expected from Eq. (15), the increase of the MSD is logarithmic and is greater the larger is $T_{com}$. Of course, we recover a true long-range order in the limit $T_{com}\rightarrow 0$ with a flat plateau of the MSD. In the inset, we extract the coefficient $\kappa$ in front of the logarithmic increase of the MSD as a function of $T_{com}$. That is:
@@ IV Conclusions and perspectives towards realistic systems
Quasi-2D vibrated granular fluids [116, 117, 118, 119, 120, 121, 122, 123, 38] represent realization of the coarse grained model discussed in section III . In these systems, macroscopic grains are confined in a box of size $L\times L\times h$ with $h\ll L$ the height of the box. The box is sinusoidally vibrated which imparts vertical energy to the grains. Collisions between particles transfer this vertical energy to the $xy$ components of the velocity. Moreover, tangential friction resulting from collisions with the top or bottom plates slows down the beads in the $xy$ plane. When viewed purely in the $xy$ plane, this system behaves similarly to the coarse-grained model discussed in section III. Indeed, collisions between particles inject energy into the $xy$ plane and play the same role as the $\Delta$ in Eq. (23) while the damping $\gamma_{com}$ during the free flight in the coarse-grained model is equivalent to the energy loss due to collisions with the roof and the bottom plate in the quasi-2D system. Moreover, the roughness of the two horizontal confining plates implies that each grain-plate collision introduces some randomness to the velocities (without even considering the angular momentum of the grains). Remarkably, this effect has been found to act as a small homogeneous thermal noise [42] which can be mapped into an effective global temperature $T_{com}$. The fact that this realistic setup is correctly described by the coarse grain model suggests the possibility of enhancing the quasi-long-range order by tuning the experimental parameters (driving amplitude, frequency, material properties, …) to reduce as much as possible the ratio between the effective $T_{com}$ caused by the plate roughness and the horizontal kinetic energy of the system $T$. Of course, the specific way in which the experimental parameters influence $T_{com}$ and $T$ can be very complex, but after some preliminary analysis, one might in principle be able to find the right combination of them to vary in order to arbitrarily reduce $T_{com}/T$. We look forward to applying these ideas to realistic granular quasi-2D systems in future studies.
@@ Appendix A Discrete real space conserving noises for vectorial quantities
The first term is a discrete Laplacian and the second is a discrete divergence of a rank 2 random tensor. Once the discrete differential operators are well defined, this equation can be understood as a discrete analogue of the model B field theory:
$m\ddot{\bm{u}}_{n}=-\gamma_{loc}\bm{\nabla}\cdot(\bm{\nabla}\bm{\dot{}}u_{n}+\sqrt{2\gamma_{loc}T_{loc}}\bm{\Xi}_{n}),$ (30)
with $\bm{\nabla}$ a discrete gradient operator. Under this form, $\bm{\Xi}$ is immediately interpreted as a flux of momentum. Note however that the global conservation of momentum depends on boundary conditions.
On a general basis, any lattice can be understood as a graph $\mathcal{G}$ with a set of vertices $\mathcal{V}$ and edges $\mathcal{E}$ connecting them. The usual notion of continuous field living on $\mathbb{R}^{d}$ (or more generally on a manifold) is replaced by the notion of field living on vertices and mapping vertex to vectors. For example in our case $\bm{u}:\mathcal{V}\to\mathbb{R}^{2}$. The complexity of dealing with discrete calculus on lattices lies in the fact that the discrete gradient maps fields living on vertices to fields living on the edges while the discrete divergence maps fields living on the edges to fields living on the vertices [124, 125]. Hence, the Laplacian being the composition of the divergence with the gradient maps fields leaving on the vertices to fields living on the vertices. This immediately implies that the divergence in Eq. (29) is ill defined if the random tensor $\bm{\Xi}$ lives on the vertices and not on the edges. Following Ref. 68, we will see that this term can be represented through a finite difference of random vertices field. Moreover, in order for the dynamics to be equilibrium-like, the discretization scheme of the differential operators as well as the variance of the noise must be well chosen in order to respect the FDT.
@@ Appendix B Fourier transform convention
Appendix C Anharmonic lattices
The theory described in Sec. II was developed with respect to an harmonic lattice. While anharmonicity leads to benign effect such as a renormalization of the harmonic spring constant, it also leads to energy transfer between modes which could change the phenomenology found for a harmonic lattice. In this section, we argue that this is not the case.
In an system with anharmonic interaction, each mode would still be driven at temperature $\tilde{T}_{k}$ as in the harmonic case, however, a transfer of energy between them would be observed. We can take as an example simple polynomial forces in real space (without loss of generality, we will stay in 1D, see Ref. 128 for a generalization in higher dimension):
$F_{\textrm{anharmonic}}=-\sum_{b=2}^{\infty}\beta_{b}\sum_{\{\bar{n}\}}(u_{n}-u_{\bar{n}})|u_{n}-u_{\bar{n}}|^{b-1},$
with $\beta_{b}$ coupling constants and $b$ the power law exponent associated to the anharmonic term. In Fourier space for the mode $k$, such force is written as [129]:
${F}_{\textrm{anharmonic}}=-\sum_{b=2}^{\infty}\beta_{b}\sum_{k_{1},k_{2},\dots,k_{b}}V^{{k},{k}_{1},{k}_{2},\dots,{k}_{b}}_{b}\times\\ \tilde{{u}}_{{k}_{1}}\tilde{{u}}_{{k}_{2}}\dots\tilde{{u}}_{{k}_{b}}\delta_{{k}+{k}_{1}+{k}_{2}+\dots+{k}_{b}}.$ (36)
The Kronecker $\delta$ ensures that the momentum is conserved by the scattering of waves. $\beta_{b}V_{b}^{k,k_{1},k_{2},\dots,k_{b}}$ is the interaction strength of a scattering due to the anharmonicity with power $b$. Such equations are common in wave turbulence [130]. For nearest neighbor interaction in 1D, we can show that:
$V_{b}^{k,k_{1},k_{2},\dots,k_{b}}\propto\sin(ka/2)\sin(k_{1}a/2)\dots\sin(k_{a}b/2),$ (37)
with $a$ the lattice spacing. These equations are making clear the fact that, with anharmonic terms, modes interact.
A proof of the absence of long-range order in systems with anharmonic interactions is beyond the scope of this work and would most likely require the full stationary distribution to be known as in the equilibrium case [2]. For example, the steady-state distribution of a three-body system, with each particle driven at a different temperature and interacting through a harmonic potential, is known, but already shows great complexity [68]. However, a qualitative argument can be made, suggesting that the overall picture presented in this section should remain largely unaffected at large length scales even in the presence of anharmonic interactions.
Since the energy is created at the level of collision $k\simeq 2\pi/a$, the most excited modes will be those around this length scale. The question that remains to be answered is: Can these modes transfer sufficiently quickly energy to the modes at $k\to 0$ so that very large wavelength phonons are excited and can reach a non-vanishing temperature? The interaction strength between these two scales is given by $V_{b}^{k\simeq 0,k_{1},\dots,k_{m}\simeq 2\pi/a,\dots,k_{b}}$ where at least one of the mode is close to 0 (long wavelength) and one mode is close to $2\pi/a$ (collisional/energy creation length scale). However, from Eq. 37, it is clear that the strength of such process goes to 0 with $k\to 0$. Hence, the rate of energy transfer to modes with very large wavelength is vanishingly small with $k$. Since the continuous damping effectively dissipates any energy supplied to these modes on a timescale by $\gamma_{com}$, the phenomenology observed for the harmonic case is expected to hold qualitatively. That is, we expect low-k mode to be roughly at temperature $\tilde{T}_{k}$, with $T_{k\to 0}=0$. Note that the argument holds in larger dimensions where the exact expression $V_{b}^{\bm{k},\bm{k}_{1},\bm{k}_{2},\dots,\bm{k}_{b}}$ contains term similar to the dispersion relation for each mode $k$, which vanish when $k\to 0$.
Appendix D Computation of the displacement correlation function
@@
Appendix E Entropons in a two temperature delta correlated bath
@@
Appendix F Hybrid event-driven/time-stepped molecular dynamics
@@
Figure 4: Effect of a global bath on the $\Delta$+$\gamma$ model for the structure factor and the translational pair correlation function. a) Main figure: structure factor as a function of the wavenumber. For lower $T_{com}$ it is easier to observe a transient hyperuniform scaling $k^{2}$. In the limit $T_{com}=0$, the hyperuniform scaling continues to $k=0$. Inset: static velocity factor divided by $T_{com}$ as a function of $k$. We obtain the limit $mC_{\dot{u}\dot{u}}(|\bm{k}|\to 0)/2=T_{com}$ as expected from $T_{k\to 0}=T_{com}$. b) Translational correlation function as a function of $r$. Increasing $T_{com}$, we observe a stronger quasi-long-range power law decay. In the limit $T_{com}=0$, $g_{\bm{G}}$ flattens and true long-range order is obtained. $N=500080$, $\phi=0.765$, $\Delta/\sigma\hat{\tau}^{-1}=0.04$, $\gamma_{com}\hat{\tau}=0.3$ and $\alpha=0.95$.
Appendix G Temperature of the active collision bath
@@
Appendix H Structure factor and translational correlation function in the presence of a global bath
In Fig. 3 we analyzed the dependence of the MSD on $T_{com}$. As described with Fig. 2 and in the theory, the plateau of the MSD, the low-$k$ structure factor and the long-range behavior of the translational pair correlation function contain the same information as given by Eqs. (14) and (17). Nonetheless, as a consistent check, it is interesting to confirm numerically these predictions in our active hard disk model. In Fig. 4a), we provide the structure factor of systems driven at different global bath temperatures. As expected, we observe that the smaller $T_{com}$ is, the lower the structure factor can go, in agreement with Eq. (13). The increase of the plateau of the MSD with system size is proportional to $S(0)$, however, in the main text we use the MSD since we consider it provides better data for finite system size (at infinite system size it would be easier to measure the plateau of the structure factor at low-$k$). In the inset, we provide as well, the velocity static factor that converges, as it should for small $k$ to $T_{k\to 0}=T_{com}$. In Fig. 4b), we give the translational pair correlation for these systems. Their long-range behavior is in agreement with Eq. (16): they exhibit a power law decay with an exponent increasing with $T_{com}$. The exponent of the power law decay is proportional to the increase of the MSD plateau, however, as for the structure factor, data are cleaner with the MSD.